# Boosting In-Silicon Directed Evolution with Fine-Tuned Protein Language Model and Tree Search

### v3

@@ Abstract
Protein evolution through amino acid mutations is a cornerstone of life sciences. Recent advances in protein language models have shown rich evolutionary patterns, offering unprecedented potential for in-silicon directed evolution. However, existing directed evolution methods largely rely on heuristic evolution strategies and have yet to efficiently integrate the transformative protein language models with advanced optimization techniques, such as reinforcement learning, to learn optimal evolution policies. To bridge this gap, we propose AlphaDE, a novel framework that evolves protein sequences by harnessing the innovative paradigms of large language models, such as fine-tuning and test-time inference. First, AlphaDE fine-tunes pretrained protein language models using masked language modeling on homologous protein sequences to activate the evolutionary plausibility of the interested protein family. Second, AlphaDE introduces test-time inference based on Monte Carlo tree search, which effectively evolves proteins with evolutionary guidance from the fine-tuned protein language model. Extensive benchmark experiments show that AlphaDE remarkably outperforms previous state-of-the-art methods even with few-shot fine-tuning. A case study further demonstrates that AlphaDE supports condensing the protein sequence space of avGFP through computational evolution.
@@ 1 Introduction
Proteins are essential components of living systems, exhibiting a great diversity of functions among biological macromolecules (holm_mapping_1996). They play critical roles in a wide range of biochemical processes, including enzyme catalysis, cellular metabolism, immune responses, and signal transduction (jiang_general_2024). Protein engineering through directed evolution enables optimization of protein functions by seeking potential protein variants with improved properties such as the expression level and catalytic activity (yang_machine-learning-guided_2019). Traditional in vitro or in vivo experimental directed evolution approaches, such as deep mutational scanning (fowler_deep_2014) and orthogonal DNA replication system (ravikumar_orthogonal_2014), directly measure the functional effects of protein mutations but are limited to exploring only a fraction of the possible protein space, and are usually expensive and laborious. To reduce the burden of the expensive wet experiments, recent advances in machine learning-guided approaches focus on building a surrogate sequence-function landscape (yang_machine-learning-guided_2019), and a lot of in-silicon directed evolution algorithms (brookes_design_2018; sinai_adalead_2020; ren_proximal_2022; wang_self-play_2023) have been proposed to strategically explore the protein fitness landscape to identify desired sequence mutations via an iterative search process.
@@
On the other hand, protein language models (meier_language_2021; lin_evolutionary-scale_2023), which encapsulate millions of years of evolutionary information through unsupervised pretraining in massive protein databases, are rapidly changing the domain of protein design (madani_large_2023; jiang_rapid_2024; hayes_simulating_2025). As protein language models implicitly learn complex evolutionary and structural dependencies from natural protein sequences, researchers have been increasingly employing protein language models for protein engineering tasks such as zero-shot inference of the functional effects of sequence substitutions to find high-fitness variations (hie_efficient_2024; shanker_unsupervised_2024). More recently, some pioneering works have begun to introduce pretrained protein language models in the process of directed evolution (jiang_rapid_2024; yang_active_2025; tran_protein_2025). However, although employing protein language models to recommend mutations, these works use simple heuristic search methods such as greedy selection (jiang_rapid_2024) or beam search (tran_protein_2025) for directed evolution. This leaves an important open question of how to seamlessly integrate protein language models into advanced optimization techniques, such as reinforcement learning, to adaptively learn optimal evolution policies.
@@
To bridge this gap, in this work, we propose a novel framework named AlphaDE to directly evolve protein sequences following the technical paradigm of natural large language models (guo_deepseek-r1_2025). Specifically, AlphaDE consists of a fine-tuning step and a Monte Carlo tree search (MCTS) inference step built on the protein language model. In the fine-tuning step, AlphaDE fine-tunes the pretrained protein language model with homologous protein sequences to activate its evolutionary plausibility for the interested protein family. In the MCTS-assisted inference step, similar to test-time MCTS in large language models to boost reasoning (zhou_language_2024; guan_rstar-math_2025), AlphaDE conducts an iterative tree search to efficiently optimize protein function via residue mutations guided by the fine-tuned protein language model. Through the two synergic steps, AlphaDE harvests superior directed evolution ability for proteins. To evaluate AlphaDE, we conduct computational benchmark experiments on eight distinct tasks. Impressively, AlphaDE substantially outperforms various in-silicon directed evolution methods. Further few-shot fine-tuning experiments reveal that AlphaDE’s evolution ability can be activated by fine-tuning with only dozens of homologous sequences. Finally, in a proof-of-concept study on avGFP, we demonstrate that AlphaDE can computationally condense the functional sequence space, highlighting its broader applicability.
@@ 2.1 Protein Language Models
In recent years, in analogy to large language models in natural language processing, protein language models such as ESM (rives_biological_2021; lin_evolutionary-scale_2023), ProteinBERT (brandes_proteinbert_2022), and ProGen (madani_large_2023) have emerged as a transformative paradigm for modeling protein sequences. Pretrained by masked language modeling (devlin_bert_2019) or autoregressive language modeling (brown_language_2020) on evolutionary-scale protein databases, protein language models have demonstrated outstanding potential in protein design. For instance, recently, hayes_simulating_2025 prompt a protein language model (i.e., ESM3) to generate a bright fluorescent protein far different from known fluorescent proteins, which is estimated to simulate five hundred million years of evolution. Another notable example is that bhat_novo_2025 utilize contrastive language modeling to design peptide binders to conformationally diverse targets using only the amino acid sequence of the target protein.
Moreover, some works (widatalla_aligning_2024; yang_steering_2025) utilize labeled sequence-fitness pairs to steer protein language models for de novo protein sequence design to generate novel and high-quality sequences, showing the superior alignment ability of protein language models.
@@ 2.2 Protein Directed Evolution
Directed evolution is a classical paradigm for protein sequence design, where plenty of algorithms are developed to accelerate the in-silicon directed evolution process. AdaLead (sinai_adalead_2020) is an advanced implementation of model-guided directed evolution with iteratively recombined and mutated operations for seed sequences. CMA-ES (hansen_completely_2001) is a second-order evolutionary search algorithm that estimates the covariance matrix to adaptively adjust the search strategy of the upcoming generations. Bayesian optimization (BO) (snoek_practical_2012) is a classical paradigm for the sequential design problem (mockus_bayesian_1989), which estimates the uncertainty and constructs the acquisition function for exploration. DbAS (brookes_design_2018) establishes a probabilistic framework that trains a variational autoencoder (VAE) (kingma_auto-encoding_2022) to model the distribution of high-fitness sequences and adaptively samples sequences from this trained VAE to explore the fitness landscape. Following DbAS, CbAS (brookes_conditioning_2019) estimates the probability distribution conditioned on the desired properties with model-based adaptive sampling and additionally considers a regularization to stabilize the model-guided search process. DyNA-PPO (angermueller_model-based_2020) formulates protein sequence design as a sequential decision-making problem and uses proximal policy optimization (PPO) (schulman_proximal_2017) to perform sequence generation. PEX (ren_proximal_2022) aims to search for effective candidates of low-order mutants near the wild-type, and formulates this process as a proximal optimization problem to solve. EvoPlay (wang_self-play_2023) uses the self-play reinforcement learning inspired by AlphaZero (silver_general_2018) to optimize protein sequences. Recently, TreeNeuralTS and TreeNeuralUCB (qiu_tree_2024) combine tree search with bandit machine learning for directed evolution, which expands a tree starting from the initial sequence with the guidance of a bandit machine learning model. Specifically, TreeNeuralTS adopts Thompson sampling (thompson_likelihood_1933) while TreeNeuralUCB utilizes the upper confidence bound (auer_finite-time_2002) to explore the sequence space.
@@
More recently, with the emergence and development of protein language models, some pioneering works explore integrating pretrained protein language models in the process of directed evolution (jiang_rapid_2024; yang_active_2025; tran_protein_2025). For example, jiang_rapid_2024 and yang_active_2025 present the multi-round active learning to rapidly improve protein activity, which alternates between collecting sequence fitness using a wet-lab assay and training a protein language model to prioritize new sequences to screen in the next round. Meanwhile, LatentDE (tran_latentde_2025) uses a gradient-based method to search in the latent space represented by VAE with protein embeddings extracted from a frozen, general-purpose pretrained protein language model (i.e., ESM-2). Moreover, MLDE (tran_protein_2025) introduces an optimization pipeline that utilizes protein language models to pinpoint the mutation hotspots and then suggest replacements by heuristically calculating a k-mer’s relevancy and entropy in a sequence. However, despite leveraging protein language models to recommend mutations, these works use simple search methods such as greedy selection (jiang_rapid_2024) or beam search (tran_protein_2025), thus leaving a critical challenge unexplored in directed evolution: how to efficiently integrate protein language models with advanced optimization techniques like reinforcement learning to further boost protein evolution. To bridge this gap, here we propose AlphaDE with a principle of effective searching (i.e., AlphaZero-like MCTS) with strong prior guidance (i.e., fine-tuned protein language model), inspired by the natural large language model technique paradigm where fine-tuning (shao_deepseekmath_2024) and test-time inference (guan_rstar-math_2025) improve problem-solving, to establish a new paradigm for in-silicon protein directed evolution.
@@ 3.1 Problem Definition
Protein directed evolution can be formulated as a Markov decision process (MDP) (bellman_markovian_1957), given that the next mutation residue to be chosen depends on the current protein sequence (wang_self-play_2023). The MDP is defined as $M=(S,A,P,R)$ where $S$ denotes the set of states that describe the current protein sequence, $A$ denotes the set of actions that indicate the chosen position and residue type to be mutated from the current protein sequence, and $P:S\times A\rightarrow S$ is the state transition function where the current protein sequence incorporates the chosen residue position and type to mutate to a new protein sequence. The protein sequence state $s\in S$ is represented by a binary matrix of size $20\times L$ with columns indicating positions ($L$) and rows indicating amino acids (20). Correspondingly, the action $a\in A$ is a flattened one-hot vector of size $20\times L$. When an action $a$ is performed, the element of value 1 in the action vector is selected, resulting in a single-site mutation on the current protein sequence $s$. The element indicated by the mutation action in the state matrix is changed to 1, while the other elements (residue types) of the same column (position in protein) are all set to 0. An episode contains a series of mutation actions $a_{t}$ and states $s_{t}$ where $0<t<T$ and $T$ is the termination step. $R:S\rightarrow\mathbb{R}$ is the episodic reward function indicating protein fitness, such as binding affinity, which can be accessed by an experimental assay or predicted by a surrogate oracle. Directed evolution aims to take actions that maximize $\overline{r}$, which is approximated under repeated rollouts (gelly_monte-carlo_2011) as
@@ 3.2 Fine-Tuning Pretrained Protein Language Model
Fine-tuning natural large language models on domain-specific data has become a standard practice to better solve specialized tasks, such as math (shao_deepseekmath_2024), coding (muennighoff2024octopack), and medical analysis (christophe2024med), by adapting large language models to enhance task-specific capabilities. Inspired by this, we fine-tune the protein language model with homologous protein sequences to activate its evolutionary plausibility for the interested protein family. Specifically, AlphaDE uses the unsupervised masked language modeling objective (devlin_bert_2019) without fitness labels. During fine-tuning, each input protein sequence is corrupted by replacing a fraction of amino acids with a special mask token. The protein language model is then trained to predict masked tokens in the corrupted sequences sampled from a homologous sequence set $S_{h}$:
@@
3.3 Monte Carlo Tree Search with Fine-Tuned Protein Language Model
We use Monte Carlo tree search (MCTS) to solve the MDP of directed evolution with guidance from protein language models, inspired by MCTS-based test-time reasoning boosts in large language models that tree-based search prompts better solutions (zhou_language_2024; guan_rstar-math_2025). Diverging from prior MCTS works that train neural networks on expert data from scratch for action guidance (rosin_multi-armed_2011; silver_mastering_2016), our approach leverages a fine-tuned protein language model to provide evolutionary guidance for mutation decisions, which predicts the next mutation action (i.e., the residue position and residue type) given the current protein sequence. Specifically, at each step, the fine-tuned protein language model receives the protein sequence and outputs the probability logits of each residue type at each residue position to guide the MCTS search.
When mutating protein sequences with the fine-tuned protein language model, although we could mutate a protein sequence in a greedy manner by taking the next mutation with the maximum probability, it is prone to being stuck in a local optimum due to the unpredictable complexity of the protein fitness landscape (romero_exploring_2009). Moreover, the predicted mutation with the maximum probability does not mean that it is optimal because the protein language model is not optimized for a specific protein fitness. To directly maximize a given protein fitness, we integrate MCTS into AlphaDE to solve the MDP of directed evolution with the help of a fine-tuned protein language model for mutation guidance. Next, we explain the MCTS part of AlphaDE in the background of protein directed evolution.
MCTS (browne_survey_2012) adopts a tree structure to perform simulation iterations and estimate the state value of actions. Meanwhile, it uses the previously estimated action values to guide the search process towards higher rewards. As shown in Figure 1, the MCTS part in AlphaDE consists of the following four steps per iteration:
  • Selection. Each iteration starts from sequence $s_{\tau}$, recursively selecting the best child node until reaching a leaf node $a_{\tau+l}$, i.e., a node that has not been expanded or terminated, after $l$ selections. At each selection step $t\in[1,l]$, a selection criterion determines the best child node to be chosen, which balances exploitation and exploration to avoid being trapped in local optimum. We use Predictor with Upper Confidence bounds applied to Trees (PUCT) (rosin_multi-armed_2011) as the selection criterion for each candidate child node as

### v4

@@ Abstract
Protein evolution through amino acid mutations is a cornerstone of life sciences. Recent advances in protein language models have shown rich evolutionary patterns, offering unprecedented potential for in-silicon directed evolution. However, existing directed evolution methods largely rely on heuristic evolution strategies and have yet to efficiently integrate the transformative protein language models with advanced optimization techniques, such as reinforcement learning, to adaptively learn superior evolution policies. To bridge this gap, we propose AlphaDE, a novel framework that evolves protein sequences by harnessing the innovative paradigms of large language models, such as fine-tuning and test-time inference. First, AlphaDE fine-tunes pretrained protein language models using masked language modeling on homologous protein sequences to activate the evolutionary plausibility of the interested protein family. Second, AlphaDE introduces test-time inference based on Monte Carlo tree search, which effectively evolves proteins with evolutionary guidance from the fine-tuned protein language model. Extensive benchmark experiments show that AlphaDE remarkably outperforms previous state-of-the-art methods even with few-shot fine-tuning. A case study further demonstrates that AlphaDE supports condensing the protein sequence space of avGFP through computational evolution.
Machine Learning, ICML
@@ 1 Introduction
Proteins are essential components of living systems, exhibiting a great diversity of functions among biological macromolecules (Holm and Sander, 1996). They play critical roles in a wide range of biochemical processes, including enzyme catalysis, cellular metabolism, immune responses, and signal transduction (Jiang et al., 2024a). Protein engineering through directed evolution enables optimization of protein functions by seeking potential protein variants with improved properties such as the expression level and catalytic activity (Yang et al., 2019). Traditional in vitro or in vivo experimental directed evolution approaches, such as deep mutational scanning (Fowler and Fields, 2014) and orthogonal DNA replication system (Ravikumar et al., 2014), directly measure the functional effects of protein mutations but are limited to exploring only a fraction of the possible protein space, and are usually expensive and laborious. To reduce the burden of the expensive wet experiments, recent advances in machine learning-guided approaches focus on building a surrogate sequence-function landscape (Yang et al., 2019), and a lot of in-silicon directed evolution algorithms (Brookes and Listgarten, 2018; Sinai et al., 2020; Ren et al., 2022; Wang et al., 2023) have been proposed to strategically explore the protein fitness landscape to identify desired sequence mutations via an iterative search process.
@@
On the other hand, protein language models (Meier et al., 2021; Lin et al., 2023), which encapsulate millions of years of evolutionary information through unsupervised pretraining in massive protein databases, are rapidly changing the domain of protein design (Madani et al., 2023; Jiang et al., 2024b; Hayes et al., 2025). As protein language models implicitly learn complex evolutionary and structural dependencies from natural protein sequences, researchers have been increasingly employing protein language models for protein engineering tasks such as zero-shot inference of the functional effects of sequence substitutions to find high-fitness variations (Hie et al., 2024; Shanker et al., 2024). More recently, some pioneering works have begun to introduce pretrained protein language models in the process of directed evolution (Jiang et al., 2024b; Yang et al., 2025b; Tran and Hy, 2025). However, although employing protein language models to recommend mutations, these works use simple heuristic search methods such as greedy selection (Jiang et al., 2024b) or beam search (Tran and Hy, 2025) for directed evolution. This leaves an important open question of how to seamlessly integrate protein language models into advanced optimization techniques, such as reinforcement learning, to adaptively learn superior evolution policies.
@@
To bridge this gap, in this work, we propose a novel framework named AlphaDE to directly evolve protein sequences following the technical paradigm of natural large language models (Guo et al., 2025). Specifically, AlphaDE consists of a fine-tuning step and a Monte Carlo tree search (MCTS) inference step built on the protein language model. In the fine-tuning step, AlphaDE fine-tunes the pretrained protein language model with homologous protein sequences to activate its evolutionary plausibility for the interested protein family. In the MCTS-assisted inference step, similar to test-time MCTS in large language models to boost reasoning (Zhou et al., 2024; Guan et al., 2025), AlphaDE conducts an iterative tree search to efficiently optimize protein function via residue mutations guided by the fine-tuned protein language model. Through the two synergic steps, AlphaDE harvests superior directed evolution ability for proteins. To evaluate AlphaDE, we conduct computational benchmark experiments on eight distinct tasks. Impressively, AlphaDE substantially outperforms various in-silicon directed evolution methods. Further few-shot fine-tuning experiments reveal that AlphaDE’s evolution ability can be activated by fine-tuning with only dozens of homologous sequences. Finally, in a proof-of-concept study on avGFP, we demonstrate that AlphaDE computationally condenses the functional sequence space, highlighting its broader applicability.
@@ 2.1 Protein Language Models
In recent years, in analogy to large language models in natural language processing, protein language models such as ESM (Rives et al., 2021; Lin et al., 2023), ProteinBERT (Brandes et al., 2022), and ProGen (Madani et al., 2023) have emerged as a transformative paradigm for modeling protein sequences. Pretrained by masked language modeling (Devlin et al., 2019) or autoregressive language modeling (Brown et al., 2020) on evolutionary-scale protein databases, protein language models have demonstrated outstanding potential in protein design. For instance, recently, Hayes et al. (2025) prompt a protein language model (i.e., ESM3) to generate a bright fluorescent protein far different from known fluorescent proteins, which is estimated to simulate five hundred million years of evolution. Another notable example is that Bhat et al. (2025) utilize contrastive language modeling to design peptide binders to conformationally diverse targets using only the amino acid sequence of the target protein.
Moreover, some works (Widatalla et al., 2024; Yang et al., 2025a) utilize labeled sequence-fitness pairs to steer protein language models for de novo protein sequence design to generate novel and high-quality sequences, showing the superior alignment ability of protein language models.
@@ 2.2 Protein Directed Evolution
Directed evolution is a classical paradigm for protein sequence design, where plenty of algorithms are developed to accelerate the in-silicon directed evolution process. AdaLead (Sinai et al., 2020) is an advanced implementation of model-guided directed evolution with iteratively recombined and mutated operations for seed sequences. CMA-ES (Hansen and Ostermeier, 2001) is a second-order evolutionary search algorithm that estimates the covariance matrix to adaptively adjust the search strategy of the upcoming generations. Bayesian optimization (BO) (Snoek et al., 2012) is a classical paradigm for the sequential design problem (Mockus, 1989), which estimates the uncertainty and constructs the acquisition function for exploration. DbAS (Brookes and Listgarten, 2018) establishes a probabilistic framework that trains a variational autoencoder (VAE) (Kingma and Welling, 2022) to model the distribution of high-fitness sequences and adaptively samples sequences from this trained VAE to explore the fitness landscape. Following DbAS, CbAS (Brookes et al., 2019) estimates the probability distribution conditioned on the desired properties with model-based adaptive sampling and additionally considers a regularization to stabilize the model-guided search process. DyNA-PPO (Angermueller et al., 2020) formulates protein sequence design as a sequential decision-making problem and uses proximal policy optimization (PPO) (Schulman et al., 2017) to perform sequence generation. PEX (Ren et al., 2022) aims to search for effective candidates of low-order mutants near the wild-type, and formulates this process as a proximal optimization problem to solve. EvoPlay (Wang et al., 2023) uses the self-play reinforcement learning inspired by AlphaZero (Silver et al., 2018) to optimize protein sequences. Recently, TreeNeuralTS and TreeNeuralUCB (Qiu et al., 2024) combine tree search with bandit machine learning for directed evolution, which expands a tree starting from the initial sequence with the guidance of a bandit machine learning model. Specifically, TreeNeuralTS adopts Thompson sampling (Thompson, 1933) while TreeNeuralUCB utilizes the upper confidence bound (Auer et al., 2002) to explore the sequence space.
@@
More recently, with the emergence and development of protein language models, some pioneering works explore integrating pretrained protein language models in the process of directed evolution (Jiang et al., 2024b; Yang et al., 2025b; Tran and Hy, 2025). For example, Jiang et al. (2024b) and Yang et al. (2025b) present the multi-round active learning to rapidly improve protein activity, which alternates between collecting sequence fitness using a wet-lab assay and training a protein language model to prioritize new sequences to screen in the next round. Meanwhile, LatentDE (Tran et al., 2025) uses a gradient-based method to search in the latent space represented by VAE with protein embeddings extracted from a frozen, general-purpose pretrained protein language model (i.e., ESM-2). Moreover, MLDE (Tran and Hy, 2025) introduces an optimization pipeline that utilizes protein language models to pinpoint the mutation hotspots and then suggest replacements by heuristically calculating a k-mer’s relevancy and entropy in a sequence. However, despite leveraging protein language models to recommend mutations, the above works use simple search methods such as greedy selection (Jiang et al., 2024b) or beam search (Tran and Hy, 2025), thus leaving a critical challenge unexplored in protein directed evolution: how to integrate protein language models with advanced optimization techniques like reinforcement learning to efficiently boost protein evolution. To bridge this gap, here we propose AlphaDE with a principle of effective searching (i.e., AlphaZero-like MCTS) with strong prior guidance (i.e., fine-tuned protein language model), inspired by the natural large language model technique paradigm where fine-tuning (Shao et al., 2024) and test-time inference (Guan et al., 2025) improve problem-solving, to establish a new paradigm for in-silicon protein directed evolution.
@@ 3.1 Problem Definition
Protein directed evolution can be formulated as a Markov decision process (MDP) (Bellman, 1957), given that the next mutation residue to be chosen depends on the current protein sequence (Wang et al., 2023). The MDP is defined as $M=(S,A,P,R)$ where $S$ denotes the set of states that describe the current protein sequence, $A$ denotes the set of actions that indicate the chosen position and residue type to be mutated from the current protein sequence, and $P:S\times A\rightarrow S$ is the state transition function where the current protein sequence incorporates the chosen residue position and type to mutate to a new protein sequence. The protein sequence state $s\in S$ is represented by a binary matrix of size $20\times L$ with columns indicating positions ($L$) and rows indicating amino acids (20). Correspondingly, the action $a\in A$ is a flattened one-hot vector of size $20\times L$. When an action $a$ is performed, the element of value 1 in the action vector is selected, resulting in a single-site mutation on the current protein sequence $s$. The element indicated by the mutation action in the state matrix is changed to 1, while the other elements (residue types) of the same column (position in protein) are all set to 0. An episode contains a series of mutation actions $a_{t}$ and states $s_{t}$ where $0<t<T$ and $T$ is the termination step. $R:S\rightarrow\mathbb{R}$ is the episodic reward function indicating protein fitness, such as binding affinity, which can be accessed by an experimental assay or predicted by a surrogate oracle. Directed evolution aims to take actions that maximize $\overline{r}$, which is approximated under repeated rollouts (Gelly and Silver, 2011) as
@@ 3.2 Fine-Tuning Pretrained Protein Language Model
Fine-tuning natural large language models on domain-specific data has become a standard practice to better solve specialized tasks, such as math (Shao et al., 2024), coding (Muennighoff et al., 2024), and medical analysis (Christophe et al., 2024), by adapting large language models to enhance task-specific capabilities. Inspired by this, we fine-tune the protein language model with homologous protein sequences to activate its evolutionary plausibility for the interested protein family. Specifically, AlphaDE uses the unsupervised masked language modeling objective (Devlin et al., 2019) without fitness labels. During fine-tuning, each input protein sequence is corrupted by replacing a fraction of amino acids with a special mask token. The protein language model is then trained to predict masked tokens in the corrupted sequences sampled from a homologous sequence set $S_{h}$:
@@
3.3 MCTS with Fine-Tuned Protein Language Model
We use Monte Carlo tree search (MCTS) to solve the MDP of directed evolution with guidance from protein language models, inspired by the test-time reasoning improvements in large language models that MCTS-based search generates better solutions (Zhou et al., 2024; Guan et al., 2025). Diverging from prior MCTS works that train neural networks on expert data from scratch for action guidance (Rosin, 2011; Silver et al., 2016), our approach leverages a fine-tuned protein language model to provide evolutionary guidance for mutation decisions, which predicts the next mutation action (i.e., the residue position and residue type) given the current protein sequence. Specifically, at each step, the fine-tuned protein language model receives the protein sequence and outputs the probability logits of each residue type at each residue position to guide the MCTS search.
When mutating protein sequences with the fine-tuned protein language model, although we could mutate a protein sequence in a greedy manner by taking the next mutation with the maximum probability, it is prone to being stuck in a local optimum due to the unpredictable complexity of the protein fitness landscape (Romero and Arnold, 2009). Moreover, the predicted mutation with the maximum probability does not mean that it is optimal because the protein language model is not optimized for a specific protein fitness. To directly maximize a given protein fitness, we integrate MCTS into AlphaDE to solve the MDP of directed evolution with the help of a fine-tuned protein language model for mutation guidance. Next, we explain the MCTS part of AlphaDE in the background of protein directed evolution.
MCTS (Browne et al., 2012) adopts a tree structure to perform simulation iterations and estimate the state value of actions. Meanwhile, it uses the previously estimated action values to guide the search process towards higher rewards. As shown in Figure 1, the MCTS part in AlphaDE consists of the following four steps per iteration:
  • Selection. Each iteration starts from sequence $s_{\tau}$, recursively selecting the best child node until reaching a leaf node $a_{\tau+l}$, i.e., a node that has not been expanded or terminated, after $l$ selections. At each selection step $t\in[1,l]$, a selection criterion determines the best child node to be chosen, which balances exploitation and exploration to avoid being trapped in local optimum. We use Predictor with Upper Confidence bounds applied to Trees (PUCT) (Rosin, 2011) as the selection criterion for each candidate child node as